Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
SASE1 Status and SASE3 Readiness Technical integration and coordination
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 651     Entry time: 24 Sep 2025, 09:43
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: Notes XO Integration Meeting 24.09.2025 
General Announcement:
- Martin Doering announces for next week's LIMP-25: 
1. PLC updates to be continued in SA3 tunnels and instruments; 
2. Karabo3 deployment starts facility wide with services being stopped; 
3. ITDM will update and replace machines.
- Timing system maintenance will from Desy will interrupt the 1.3GHz RF main refence signal tomorrow, 25.09; in October a full maintenance will be done during the first two weeks. 
Date(s) to be announced by XO when know. In case of date restrictions from EuXFEL groups, please contact XO.

CTRL:
- Nothing major to report. Tomorrow timing is required and may be provided by EEE, to be discussed offline.

DA:
- Nothing to report; calibration pipelines are being updated during the LIMP-25.

EEE:
- Nothing to report.

VAC:
- Not present.

XPD:
- Activities are ongoing.

DET:
- Tomorrow Karabo3 with AGIPD is planned to be tested; thus timing needs to be simulated (Train IDs) if not available.

ITDM: 
- Nothing to report.

XRO:
- Not present.

LAS:
- Not present.

UNSYS:
- Not present.

SPB:
- Nothing to report.

FXE:
- Nothing to report.

MID:
- Nothing to report.

HED:
- Not present.

SCS:
- Not present / nothing to report.

SQS:
- Not present.

SXP:
- Nothing to report.
ELOG V3.1.4-7c3fd00